For the equation y = 3x + 5 write the equation of the lines in slope intercept form that pass through point (-4,1) and are parallel and perpendicular to the given line.
Answer by mananth(16949)

1 y = 3 x + 5
Divide by 1
y = 3 x + 5
Compare this equation with y=mx+b
slope m = 3
The slope of a line parallel to the above line will be the same
The slope of the required line will be 3
m= 3 ,point ( -4 , 1 )
Find b by plugging the values of m & the point in
y=mx+b
1 = -12 + b
b= 13
m= 3
Plug value of the slope and b in y = mx +b
The required equation isy = 3 x + 13
For perpendicular line the slope will be negative reciprocal.
which is -1/3
the line passes through (-4,1)
y= mx +b
1=-1/3(-4)+b
1=4/3+b
b=-1/3
The equation will be y= -(1/3) x -1/3


m.ananth@hotmail.ca